OpenStudy (anonymous):

What is the simplified form of 1 over x minus 2 over x squared minus x ? (1/x ) - (2/x^2-x) Pleaseeee help These are the choices x minus 1 over the quantity x times x plus 1 1 minus x over the quantity x times x plus 1 3 minus x over the quantity x times x minus 1 x minus 3 over the quantity x times x minus 1

OpenStudy (anonymous):

I think it is the third choice.

OpenStudy (xapproachesinfinity):

\(\huge \frac{1}{x}-\frac{2}{x^2-x}=\frac{1}{x}-\frac{2}{x(x-1)}\) \(\huge \frac{\color{red}{(x-1)}}{x\color{red}{(x-1)}}-\frac{2}{x(x-1)}\) \(\huge \frac{(x-1)-2}{x(x-1)}\)

OpenStudy (xapproachesinfinity):

finish the work...

OpenStudy (xapproachesinfinity):

any problem understanding what i did? just tell me i didn't get the line number two or line number 1 i will explain